HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ROTATIONAL SPECTROSCOPY OF^{13}C N-PROPYL CYANIDE USING A HOME-BUILT MM WAVE CHIRPED-PULSE SPECTROMETER WITH A

NEW DIGITIZER

\underline{B. HEYNE}, P. MISRA, I. Physikalisches Institut, Universität zu Köln, Zülpicher Straße 77, 50937 Köln, Germany; G. GRUTZECK, Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, 53121 Bonn, Germany; M. FATIMA, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Straße

77, 50937 Köln, Germany

So far, about 300 molecules have been found in the interstellar and circumstellar medium^{1}. Among them is n-propyl cyanide, which was detected towards Sagittarius B2(N) in Atacama Large Millimeter/Submillimeter Array (ALMA) data^{2}^{,}^{3}. Since the ratio of ^{12}C/^{13}C isotopologues is close to 25 for many molecules in Sagittarius B2(N)^{4}, it is reasonable to expect that its ^{13}C isotopologues can also be observed in this source.

We use an upgrade of our home-built mm wave chirped-pulse spectrometer in the range between 75-110 GHz^{5} to measure spectra of the ^{13}C isotopologues of npropyl cyanide for the first time. The isotopologues are present in the sample in natural abundance (ca. 1%). The detector of our spectrometer employs a heterodyne receiver which was originally built for the use in an emission spectrometer^{6}. This device was not equipped for sideband separation. This makes it difficult to assign individual rotational transitions, as mirror lines are created, meaning artificial lines in the spectrum appear. Those features arising from intense lines are hard to distinguish from weak lines associated with the isotopologues. Therefore, we incorporated a new backend spectrometer into our setup, which consists of a home-built FPGA-ADC system, where we can record free induction decays for several different excitation phases in a cyclic manner to perform a sideband separation.

The new backend spectrometer in our mm wave chirped-pulse spectrometer is explained, its special features are discussed and spectra of n-propyl cyanide in natural abundance are presented which demonstrates the sensitivity of the improved instrument.

1The Cologne Database for Molecular Spectroscopy, Access on 12.06.2023.

2A. Belloche, R. T. Garrod, H. S. P. Müller, K. M. Menten, C. Comito and P. Schilke, A&A, Volume 499, 215 - 232 (2009).

3H. S. P. Müller, A. Walters, N. Wehres, A. Belloche, O. H. Wilkins, D. Liu, R. Vicente, R. T.

Garrod, K. M. Menten, F. Lewen and S. Schlemmer, A&A 595, A87 (2016).

4H. S. P. Müller, A. Walters, N. Wehres, A. Belloche, O. H. Wilkins, D. Liu, R. Vicente, R. T.

Garrod, K. M. Menten, F. Lewen and S. Schlemmer, A&A 595, A87 (2016).

5M. Hermanns, N. Wehres, B. Heyne, C. E. Honingh, U. U. Graf, S. Schlemmer, Review of Scientific Instruments 94.3, 034705 (2023).

6N. Wehres, B. Heyne, F. Lewen, M. Hermanns, B. Schmidt, C. Endres, U. U. Graf, D. R. Higgins and S. Schlemmer, Proc. Int. Astron. Union, 332-345, (2017).

p-number: p097  Submitted on Tue, 13 Jun 2023 14:43:08 +0200